# Algorithms for the Shortest Vector Problem in $2$-dimensional Lattices, Revisited

### v1

@@ V Optimized algorithm and its analysis
For two integers with bit lengths of $n$, the well-known Half-GCD algorithm [9, 10] can efficiently find two integers of approximately $n/2$-bit length that share the same common divisor. Inspired by the Half-GCD algorithm, this section first explores the HVec algorithm, which rapidly reduces a lattice basis to a new basis with bit lengths approximately halved from the original, along with a complexity analysis. Furthermore, we leverage this algorithm to optimize and accelerate the CrossEuc-SBP algorithm presented in the previous section.
@@ V-A HVec algorithm and its analysis
V-B HVecSBP algorithm and its analysis
@@ VI-A Experimental Methodology
It is evident that the practical performance of existing algorithms is closely related to the form of the input lattice basis. Therefore, we separately compare the performance for lattice bases in the special HNF and those in general form.
(1) Input lattice basis in HNF: Consider the input lattice basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a&b\\ 0&c\end{pmatrix}$, where $a>b>|c|>0$. Note that in this case, our algorithm CrossEuc is identical to the continued fraction-based algorithm (ParEuc) [4, 3]. Thus, we only compare three algorithms: the optimized Lagrange algorithm under the $\ell_{\infty}$ metric (GolEuc) [5, 3], the continued fraction-based algorithm (ParEuc) [4, 3], and our optimized algorithm (HVec-SBP).
(2) Input lattice basis in general form: For the input lattice basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}$, we evaluate the practical performance of each algorithm based on the linear dependency between vectors $\mathbf{a}$ and $\mathbf{b}$. We define the linear dependency measure $\Delta(\mathbf{a},\mathbf{b})$ as:
@@
It is clear that a smaller $\Delta$ value indicates a higher degree of linear dependency between vectors $\mathbf{a}$ and $\mathbf{b}$. In this case, we assessed the practical performance of the following six algorithms across different $\Delta$ values:
@@
  • EEA-HNF-ParEuc: An algorithm that first transforms the basis into Hermite Normal Form (HNF) using the extended Euclidean algorithm (EEA) and then reduces it with the continued fraction-based algorithm (ParEuc). Visually
@@ VI-B Experimental Environment and Result Analysis
(1) For the input lattice basis in HNF, let #$\#a=\#b=n_{1},\#c=n_{2}$, TABLE I compares the time cost of each algorithms as the variance of $n_{2}$. These results highlight HVecSBP’s superior performance in handling HNF-structured bases, particularly in scenarios with highly imbalanced parameter lengths ($n_{2}\ll n_{1}$). Such scenarios are common in cryptographic applications, such as lattice-based attacks, where HVecSBP achieves at least 13.5×13.5$\times$13.5 × efficiency improvement over previous approaches.
TABLE I: Time cost of different algorithms as $n_{2}$ varies (in seconds)
@@
(2) For the input lattice basis in general form, Table II compares the time cost of each algorithm as the difference $\Delta=|a_{1}/b_{1}-a_{2}/b_{2}|\approx 10^{-\delta}$ varies. The last column, $\kappa$, represents the proportion of differing terms in the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$, counted from right to left, relative to the total number of terms in the expansions. Specially, the first $\kappa=1$ indicates that only the last term of the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differs. Notably, as $\Delta$ decreases, $\kappa$ increases, indicating that $a_{1}/b_{1}$ and $a_{2}/b_{2}$ become closer to being equal. This further implies that the lattice basis vectors $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$ approach linear dependence. From the experimental results, the following observations can be made:
@@
  • For the three algorithms without HGCD optimization—CrossEuc, EEA-HNF-ParEuc, and GolEuc: In general, the closer the two input lattice basis vectors are to being linearly dependent, the higher the time cost for these three algorithms. However, our proposed algorithm, CrossEuc, significantly outperforms the other two. Specifically, the time cost of CrossEuc is approximately $20\%$ of that of GolEuc, representing a 5×5$\times$5 × efficiency improvement. Compared to EEA-HNF-ParEuc, CrossEuc achieves a time cost of $22\%-5\%$ as $\Delta$ increases, corresponding to an efficiency improvement of 4.6×4.6$\times$4.6 × to 19×19$\times$19 ×.
  • For the HGCD-optimized algorithms—HGCD-HNF-ParEuc, HGCD-HNF-HVecSBP, and our algorithm HVecSBP, we can observe the following two experimental phenomena:
    1. 1.
@@
      When the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differ only in the last term, HGCD-HNF-ParEuc achieves the highest efficiency, followed by HGCD-HNF-HVecSBP. The time cost of HVecSBP is approximately 2×2$\times$2 × that of the other two algorithms, making it the least efficient. This is intuitive, as the input lattice basis vectors nearly degenerate into two integers, making integer-based HGCD processing more efficient. The vectorized HVec algorithm, which handles two dimensions, naturally incurs double the time cost.
    2. 2.
      As the linear dependency between the two lattice basis vectors weakens—specifically, when the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differ in approximately the last $27.5\%$ of terms—our algorithm HVecSBP begins to outperform the other two. As this proportion increases, the efficiency advantage of HVecSBP grows progressively.
TABLE II: Time cost of different algorithms as $\delta$ varies (in seconds):#$\#\{a_{1},a_{2},b_{1},b_{2}\}=10^{6},\Delta\approx 10^{-\delta}$

### v2

@@ IV-B Algorithm Description and Analysis
Remark 1.
According to the property established in Theorem 2, the $\ell_{2}$-shortest basis $[\mathbf{u}\ \mathbf{v}]$ can be obtained via a minor modification of Algorithm 3. Specifically, by replacing Steps 1114 with the following:
$\mathbf{u}=\arg\min\left\{\|\mathbf{a}\|_{2},\ \|\mathbf{b}\|_{2},\ \|\mathbf{a}+\mathbf{b}\|_{2},\ \|\mathbf{a}-\mathbf{b}\|_{2}\right\},$
and setting
$\mathbf{v}=\mathbf{x}-q\mathbf{u},\quad\text{where }\mathbf{x}\in\{\mathbf{a},\mathbf{b}\}\setminus\{\mathbf{u}\}\text{ and }q=\left\lceil\frac{\langle\mathbf{x},\mathbf{u}\rangle}{\langle\mathbf{u},\mathbf{u}\rangle}\right\rfloor.$
This substitution yields a basis satisfying the $\ell_{2}$-shortest condition. The same remark is also applicable to the upcoming Algorithm 5 (HVecSBP).
@@ V Optimized algorithm and its analysis
For two integers with bit lengths of $n$, the well-known Half-GCD algorithm [10, 11] can efficiently find two integers of approximately $n/2$-bit length that share the same common divisor. Inspired by the Half-GCD algorithm, this section first explores the HVec algorithm, which rapidly reduces a lattice basis to a new basis with bit lengths approximately halved from the original, along with a complexity analysis. Furthermore, we leverage this algorithm to optimize and accelerate the CrossEuc algorithm presented in the previous section.
@@ V-A HVec algorithm and its analysis
V-B HVecSBP algorithm and its complexity analysis
@@ VI-A Experimental Methodology
To comprehensively evaluate the practical efficiency of our proposed algorithm CrossEuc and its optimized variant HVecSBP against existing methods, we will conduct extensive experiments considering both metric types and lattice basis forms. In terms of the metric, we compare the practical performance of various algorithms under both the Euclidean and Minkowski metrics. Regarding the input basis, it is evident that the performance of existing algorithms is highly dependent on the structure of the input lattice basis. We focus on two common types: the HNF-form basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a&b\\ 0&c\end{pmatrix}$, where $a>b>|c|>0$, and the general-form basis $\mathbf{B}=[\mathbf{a}\ \mathbf{b}]=\begin{pmatrix}a_{1}&b_{1}\\ a_{2}&b_{2}\end{pmatrix}$. For general-form bases, the degree of linear dependence between the two basis vectors clearly affects algorithmic performance. We define the linear dependency measure $\Delta(\mathbf{a},\mathbf{b})$ as:
@@
It is clear that a smaller $\Delta$ value indicates a higher degree of linear dependency between vectors $\mathbf{a}$ and $\mathbf{b}$.
(1) Evaluation for $\ell_{2}$-shortest basis. We compare five algorithms: our proposed fundamental reduction algorithm CrossEuc, Lagrange reduction algorithm LagRed under the $\ell_{2}$ metric [6, 9] (see Appendix B), Yap’s fundamental algorithm CRS [15] (see Appendix C), our optimized algorithm HVecSBP and Yap’s optimized algorithm (HalfGaussianSBP) [15] (see Appendix D).
  • For HNF-form bases, we assessed the practical performance of the above-mentioned algorithms across different bit sizes of $c$.
  • For general-form bases, we assess the practical performance of the above-mentioned algorithms across different $\Delta$ values.
(2) Evaluation for $\ell_{\infty}$-shortest basis.
  • For HNF-form bases, our algorithm CrossEuc is identical to the continued fraction-based algorithm (ParEuc) [4, 3]. Thus, we only compare three algorithms: the optimized Lagrange algorithm under the $\ell_{\infty}$ metric (GolEuc) [6, 3], the continued fraction-based algorithm (ParEuc) [4, 3], and our optimized algorithm (HVecSBP).
  • For general-form bases, we evaluate the practical performance of the following algorithms across different $\Delta$ values.
@@
    • EEA-HNF-ParEuc: An algorithm that first transforms the basis into HNF using the extended Euclidean algorithm (EEA) and then reduces it with the continued fraction-based algorithm (ParEuc). Visually
@@ VI-B Experimental Environment and Result Analysis
VI-B1 Experimental result and analysis for the $\ell_{2}$-shortest basis
  • For the input lattice basis in Hermite Normal Form (HNF), Table I compares the running time of each algorithm as $n_{2}$ varies, where $n_{1}=2\times 10^{5}$ denotes the number of decimal digits in $a$ and $b$, and $n_{2}$ represents the number of decimal digits in $c$. As shown in Table I, smaller values of $n_{2}$ lead to higher computational times for all algorithms. In addition, the following observations can be made:
    • (i)
      For the three fundamental reduction algorithms CrossEuc, CRS, and LagRed, our proposed CrossEuc significantly outperforms the others, with time consumption approximately only 0.19% of CRS and 0.28% of LagRed, achieving speedups of around 500×500$\times$500 × and 350×350$\times$350 ×, respectively.
    • (ii)
      For the two optimized algorithms HalfGaussianSBP and HVecSBP, our optimized algorithm HVecSBP takes only about $7\%$ of the time used by Yap’s HalfGaussianSBP, resulting in a speedup of approximately 14×14$\times$14 ×.
    • (iii)
      For our proposed fundamental reduction algorithm $\mathbf{CrossEuc}$ and its optimized variant $\mathbf{HVecSBP}$, the runtime of HVecSBP constitutes $18.8\%$-$38.1\%$ of CrossEuc’s as $n_{2}$ decreases – achieving a $5.3$ to $2.6$-fold efficiency gain. Remarkably, $\mathbf{CrossEuc}$ itself outperforms Yap’s $\mathbf{HalfGaussianSBP}$, requiring only $35.6\%$-$18.9\%$ of its runtime under diminishing $n_{2}$. This advantage is particularly pronounced for smaller-scale lattice bases (characterized by parameter $n_{1}$).
    TABLE I: Running time of various algorithms for the $\ell_{2}$-shortest basis with varying $n_{2}$ (in seconds)
    length (dec) $n_{1}$  length (dec) $n_{2}$  CrossEuc  CRS  LagRed  HalfGaussianSBP  HVecSBP
    200000 200000 0.000522 0.035561 0.034960 0.072230 0.001405
    200000 180000 0.215646 112.676702 77.524088 0.605316 0.040622
    200000 160000 0.394690 213.510331 146.752125 1.156759 0.078501
    200000 140000 0.557541 296.380581 204.626063 1.689438 0.109761
    200000 120000 0.694540 366.024648 255.231900 2.243587 0.152746
    200000 100000 0.803900 423.430425 296.717711 2.769396 0.197548
    200000 80000 0.902036 464.113459 323.928026 3.277750 0.222285
    200000 60000 0.952363 498.170153 346.013871 3.902740 0.262041
    200000 40000 1.014240 517.602416 358.009027 4.380847 0.312382
    200000 20000 1.037861 529.151738 365.802161 4.944157 0.350908
    200000 10000 1.044446 530.145559 370.650970 5.307488 0.366273
    200000 1000 1.046709 531.093756 371.334628 5.534359 0.394135
    200000 100 1.052189 531.633906 371.504885 5.561485 0.394579
    200000 1 1.055789 531.637407 371.524263 5.567000 0.402265
  • For the input lattice basis in general form, Table II compares the time cost of each algorithm for a fixed size $n=2\times 10^{5}$, as the difference $\Delta=|a_{1}/b_{1}-a_{2}/b_{2}|\approx 10^{-\delta}$ varies, where $n$ denotes the maximum number of decimal digits in $a_{1},a_{2},b_{1}$ and $b_{2}$. The last column, $\kappa$, represents the proportion of differing terms in the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$, counted from right to left, relative to the total number of terms in the expansions. Specially, the first $\kappa=1$ indicates that only the last term of the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differs. Notably, as $\Delta$ decreases, $\delta$ increases and $\kappa$ decreases, indicating that $a_{1}/b_{1}$ and $a_{2}/b_{2}$ become closer to being equal. This further implies that the lattice basis vectors $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$ approach linear dependence. As shown in Table II, the closer the two input lattice basis vectors are to being linearly dependent ($\delta$ increases), the higher the computational time required for all algorithms. Additionally, the following observations can be made:
    • (i)
[…]
    • (ii)
[…]
    • (iii)
      For our proposed fundamental reduction algorithm $\mathbf{CrossEuc}$ and its optimized variant $\mathbf{HVecSBP}$, the runtime of the latter constitutes approximately $39.5\%$-$19.9\%$ of the former as the linear dependence between input lattice basis vectors decreases – achieving a $2.5$ to $5$-fold efficiency gain. Strikingly, $\mathbf{CrossEuc}$ itself outperforms Yap’s optimized algorithm $\mathbf{HalfGaussianSBP}$, requiring only $14.9\%$-$36.4\%$ of the latter’s runtime under diminishing linear dependence. This advantage is particularly pronounced for smaller-scale lattice bases (characterized by parameter $n$).
TABLE II: Running time of various algorithms for the $\ell_{2}$-shortest basis with varying $\delta$ (in seconds)
$\delta$  CrossEuc   CRS  LagRed  HalfGaussianSBP  HVecSBP  $\kappa$
400000 2.187857 1061.482896 732.066229 14.621288 0.865833 1
350000 2.039634 1029.768712 675.253707 9.911299 0.705745 12.5%
300000 1.963530 953.469648 640.119371 8.779507 0.597472 25%
250000 1.760551 890.452219 595.714993 7.692156 0.530357 37.5%
200000 1.541461 820.747442 564.586658 6.102788 0.430760 50%
150000 1.361687 705.880787 461.356443 4.899072 0.344573 62.5%
100000 1.059967 573.571627 395.397606 3.676309 0.260802 75%
50000 0.753551 416.365192 283.668019 2.415546 0.168508 87.5%
20000 0.571839 306.532068 206.223716 1.718590 0.131812 95%
10000 0.488057 263.113941 178.243286 1.572602 0.103363 97.5%
2000 0.458032 227.192431 159.118865 1.255080 0.090067 99.5%
1000 0.446958 224.975630 158.739315 1.243584 0.087849 99.75%
200 0.439207 223.499143 153.864744 1.234583 0.085077 99.95%
100 0.432651 222.870675 153.261555 1.233634 0.084719 99.975%
20 0.423252 222.065486 152.929364 1.232835 0.084328 99.995%
0 0.422037 221.981643 152.823700 1.231388 0.084183 100%
VI-B2 Experimental result and analysis for the $\ell_{\infty}$-shortest basis
@@
  • For the input lattice basis in HNF, TABLE III compares the time cost of each algorithms as the variance of $n_{2}$, where $n_{1}=10^{6}$ denotes the number of decimal digits in $a,b$, and $n_{2}$ represents the number of decimal digits in $c$. These results highlight HVecSBP’s superior performance in handling HNF-structured bases, particularly in scenarios with highly imbalanced parameter lengths ($n_{2}\ll n_{1}$). Such scenarios are common in cryptographic applications, such as lattice-based attacks, where HVecSBP achieves at least 13.5×13.5$\times$13.5 × efficiency improvement over previous approaches.
    TABLE III: Running time of various algorithms for the $\ell_{\infty}$-shortest basis with varying $n_{2}$ (in seconds)
@@
  • For the input lattice basis in general form, Table IV compares the time cost of each algorithm for a fixed size $n=10^{6}$, as the difference $\Delta=|a_{1}/b_{1}-a_{2}/b_{2}|\approx 10^{-\delta}$ varies, where $n$ denotes the maximum number of decimal digits in $a_{1},a_{2},b_{1}$ and $b_{2}$. The meaning of the parameter $\kappa$ is same as that in Table II. Notably, as $\Delta$ decreases, $\delta$ increases and $\kappa$ decreases, suggesting that $\frac{a_{1}}{b_{1}}$ and $\frac{a_{2}}{b_{2}}$ become increasingly similar. This, in turn, implies that the lattice basis vectors $\mathbf{a}=(a_{1}\ a_{2})^{T}$ and $\mathbf{b}=(b_{1}\ b_{2})^{T}$ are approaching linear dependence. As shown in Table IV, the following observations can be made:
    • (i)
@@
      For the three fundamental algorithms without HGCD optimization—CrossEuc, EEA-HNF-ParEuc, and GolEuc—the closer the two input lattice basis vectors are to being linearly dependent, the more computational time is required by all three. Meanwhile, our proposed algorithm, CrossEuc, consistently outperforms the other two. Compared to GolEuc, CrossEuc demonstrates a relatively stable efficiency advantage that is largely unaffected by the linear dependency of the input lattice basis vectors. Specifically, the time cost of CrossEuc is approximately 20% of that of GolEuc, representing a 5× speedup. In contrast, when compared to EEA-HNF-ParEuc, the efficiency gain of CrossEuc is highly sensitive to the linear dependency of the input vectors. As the degree of dependency decreases (i.e., as $\delta$ decreases), CrossEuc achieves a speedup ranging from 4×4$\times$4 × to 18×18$\times$18 ×, with the advantage becoming increasingly significant.
    • (ii)
@@
      For the HGCD-optimized algorithms—HGCD-HNF-ParEuc, HGCD-HNF-HVecSBP, and our algorithm HVecSBP, we have (1) When the continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ differ only in the last term, HGCD-HNF-ParEuc achieves the highest efficiency, followed by HGCD-HNF-HVecSBP. The time cost of HVecSBP is approximately 2×2$\times$2 × that of the other two algorithms, making it the least efficient. This is intuitive, as the input lattice basis vectors nearly degenerate into two integers, making integer-based HGCD processing more efficient. The vectorized HVec algorithm, which handles two dimensions, naturally incurs double the time cost. (2) As linear dependence between lattice basis vectors diminishes ($\delta$ decreases)—specifically when continued fraction expansions of $a_{1}/b_{1}$ and $a_{2}/b_{2}$ diverge in approximately the last 27.5% of terms—our algorithm $\mathbf{HVecSBP}$ achieves dominant performance. Beyond this threshold, its efficiency advantage scales progressively: accelerating from marginal parity 1×1$\times$1 × to 13-fold speedup over $\mathbf{HGCD\text{-}HNF\text{-}HVecSBP}$, and from 2×2$\times$2 × to 179×179$\times$179 × acceleration against $\mathbf{HGCD\text{-}HNF\text{-}ParEuc}$, with gains intensifying at smaller $\delta$ values.
    • (iii)
      For our proposed fundamental reduction algorithm, $\mathbf{CrossEuc}$, and its optimized variant, $\mathbf{HVecSBP}$, the runtime of $\mathbf{HVecSBP}$ is approximately 4.2% to 8.5% of that of $\mathbf{CrossEuc}$ as the linear dependence between the input lattice basis vectors decreases—resulting in an efficiency gain of roughly $\times1$ to 23×23$\times$23 ×.
TABLE IV: Running time of various algorithms for the $\ell_{\infty}$-shortest basis with varying $\delta$ varies (in seconds)